The initial temperature of a furnace is measured as 2450.00k.the temperature of furnace rises to 1/1*10 times of its initial value after an hour .what will be the increased temperature of the furnace n degree celsus

1
Expert's answer
2022-04-28T01:54:08-0400

2450.00 k=2450-273=2177 C

Increasing in 1/1*10=10 times is 2177x10=21770 C